How many gallons of a 12% salt solution must be mixed with 6 gallons of a 20% salt solution to obtain 13% salt solution?
Answer by mananth(16946)

percent percentage ---------------- quantity
Salt solution 12 ----------------x gallons
Salt solution 20 ----------------6-x gallons
Mixture 13.00% ---------------- 6

12x+20(6-x)=13* 6
12x+120-20x=78
12x-20x =78-120
-8x=-42
/-8
x=5.25 gallons 12.00% Salt solution
0.75 gallons 20.00% Salt solution